# High-order WENO finite-difference methods for hyperbolic nonconservative systems of Partial Differential Equations

### v1

@@ 4.1 Homogeneous problems
We consider uniform Cartesian meshes with points $(x_{i},y_{j})$ with steps $\Delta x$ and $\Delta y$ in the $x$ and $y$ direction. WENO methods can be extended dimension by dimension:
@@
$D^{\pm}_{i,k;j}$ $=$ $\int_{0}^{1}A_{1}^{\pm}(\Psi(s;U_{i,j},U_{k,j}))\partial_{s}\Psi(s;U_{i,j},U_{k,j})\,ds,$ (42)
$D^{\pm}_{i;j,k}$ $=$ $\int_{0}^{1}A_{2}^{\pm}(\Psi(s;U_{i,j},U_{i,k}))\partial_{s}\Psi(s;U_{i,j},U_{i,k})ds,$ (43)
where $A_{j}^{\pm}(U)$ represents the splitting chosen for $A_{j}(U)$: LF splitting can be readily extended to 2D and the upwind splitting as well if a Roe linearization is available. The extension of the strategy described in Section 2.1 to reconstruct characteristic variables is straightforward as well.
Proposition 1 can be easily extended to prove that (37) is a numerical method of order $p=2k+1$.
@@ 4.2 Problems with source terms
where $H(x,y)$ is again a known function. When a Roe linearization is available, Strategy 1 described in Section 3 can be readily extended to problem (44): it is enough to redefine the fluctuations as follows:
$D^{\pm}_{i,k;j}$ $=$ $A^{\pm}_{1;i,k;j}(U_{k,j}-U_{i,j})-S^{\pm}_{1;i,k;j}(H(x_{k},y_{j})-H(x_{i},y_{j})),$ (45)
$D^{\pm}_{i;j,k}$ $=$ $A^{\pm}_{2;i;j,k}(U_{i,k}-U_{i,j})-S^{\pm}_{2;i;j,k}(H(x_{i},y_{k})-H(x_{i},y_{j})),$ (46)
where $A_{1;i,k;j}$, $S_{1;i,k;j}$, represent respectively the intermediate matrix and source term given by the Roe linearization for the states $U_{i,j}$ and $U_{k,j}$ and $A_{2;i;j,k}$, $S_{2;i;j,k}$ the corresponding to the states $U_{i,j}$, $U_{i,k}$.
If (29) is assumed, it can be shown as in Proposition 2 that the numerical method is well-balanced for stationary solutions that satisfy that given $(x_{L},y),(x_{R},y),(x,y_{D}),(x,y_{U})\in\mathbb{R}^{2}$ the functions
@@
$s\in[0,1]\to\Psi_{U}\left(s;\left[\begin{array}[]{c}U^{*}(x,y_{D})\\ H(x,y_{U})\end{array}\right];\left[\begin{array}[]{c}U^{*}(x,y_{D})\\ H(x,y_{U})\end{array}\right]\right),$ (48)
@@
Strategy 2 described in Section 3.2 cannot be extended in general since, to implement it, a stationary solution $U^{*}_{i,j}$ satisfying
$U^{*}_{i,j}(x_{i},y_{j})=U_{i,j}$
is necessary, but now stationary solutions satisfy the PDE system
$A_{1}(U^{*})\partial_{x}U^{*}+A_{2}(U^{*})\partial_{y}U^{*}=S_{1}(U^{*})H_{x}+S_{2}(U^{*})H_{y},$
and one point value is not enough to determine one of them: in general, there will be infinitely many or none. Nevertheless, Strategy 2 can be extended easily to 2D problems to design numerical methods that preserve a given family of $m$-parameter stationary solutions
@@
with $m<N$. As in the 1D case, first the value of the parameters are imposed so that the value of $m$ pre-determined components at $(x_{i},y_{j})$ of the corresponding element of the family coincide with those of $U_{i,j}$. Once an element of the family $U^{*}_{i,j}$ has been determined, the numerical solution is written as follows:
@@ 5 Application to the two-layer shallow water model
In this Section, we consider the 2D two-layer shallow water system that governs the flow of two superposed layers of immiscible fluids with different constant densities:
@@
$U=\left[\begin{array}[]{c}h_{1}\\ h_{1}u_{1,1}\\ h_{1}u_{1,2}\\ h_{2}\\ h_{2}u_{2,1}\\ h_{2}u_{2,2}\end{array}\right],$
$A_{1}(U)=\left[\begin{array}[]{cccccc}0&1&0&0&0&0\\ c_{1}^{2}-u_{1,1}^{2}&2u_{1,1}&0&c_{1}^{2}&0&0\\ -u_{1,1}u_{1,2}&u_{1,2}&u_{1,1}&0&0&0\\ 0&0&0&0&1&0\\ rc_{2}^{2}&0&0&c_{2}^{2}-u_{2,1}^{2}&2u_{2,1}&0\\ 0&0&0&-u_{2,1}u_{2,2}&u_{2,2}&u_{2,1}\end{array}\right],$ (57)
$A_{2}(U)=\left[\begin{array}[]{cccccc}0&0&1&0&0&0\\ -u_{1,1}u_{1,2}&u_{1,2}&u_{1,1}&0&0&0\\ c_{1}^{2}-u_{1,2}^{2}&0&2u_{1,2}&c_{1}^{2}&0&0\\ 0&0&0&0&0&1\\ 0&0&0&-u_{2,1}u_{2,2}&u_{2,2}&u_{2,1}\\ rc_{2}^{2}&0&0&c_{2}^{2}-u_{2,2}^{2}&0&2u_{2,2}\end{array}\right],$ (58)
$S_{1}(U)=\left[\begin{array}[]{c}0\\ gh_{1}\\ gh_{1}\\ 0\\ 0\\ 0\end{array}\right],\quad S_{2}(U)=\left[\begin{array}[]{c}0\\ 0\\ 0\\ 0\\ gh_{2}\\ gh_{2}\end{array}\right].$ (59)
@@
The family of straight segments will be considered here to compute non-conservative products:
$\Psi(s;U_{L},U_{R})=U_{L}+s(U_{R}-U_{L}),\quad s\in[0,1].$
Steady-states solutions that correspond to water-at-rest equilibria constitute a 2-parameter family:
$u_{1,1}=u_{1,2}=u_{2,1}=u_{2,2}\equiv 0,\;h_{1}=C_{1},\;h_{2}=-Z+C_{2},$
where $C_{1}$, $C_{2}$ are arbitrary parameters ($C_{1}>0$, $C_{2}>\max(Z)$). We will focus here on methods that preserve this family.
@@
where now $u_{k}$, $h_{k}$, $k=1,2$ are respectively the velocity and thickness of the layers.
The following Roe linearization corresponding to the family of straight segments is available (see [25]):
$A(U_{L},U_{R})=\left[\begin{array}[]{ccccc}0&1&0&0\\ -\bar{u}_{1}^{2}+\bar{c}_{1}^{2}&2\bar{u}_{1}&\bar{c}_{1}^{2}&0\\ 0&0&0&1\\ r\bar{c}_{2}^{2}&0&-\bar{u}_{2}^{2}+\bar{c}_{2}^{2}&2\bar{u}_{2}\end{array}\right],\quad S(U,V)=\left[\begin{array}[]{c}0\\ g\bar{h}_{1}\\ 0\\ g\bar{h}_{2}\end{array}\right],$
@@
This Roe matrix and its natural extension to 2D will be used to implement WENO methods.
Both Strategies 1 and 2 described in Section 3 for 1D problems and 4.2 for 2D problems can be followed to design numerical methods that preserve water-at-rest solutions. In effect, for Strategy 1, the equality of the curves given by (47) and (49) or those given by (48) and (50) can be easily checked if the family of straight segments is chosen: it derives from the linear nature of the relationships between variables that characterize water-at-rest solutions. On the other hand, if Strategy 2 is followed, $U^{*}_{i,j}$ is chosen as follows;
$U^{*}_{i,j}(x,y)=\left[h_{1;i,j},0,0,-Z(x,y)+Z(x_{i},y_{j})+h_{2;i,j},0,0\right]^{T}.$
When the upwind-based splitting is used, systems of the form
$R\cdot\vec{\alpha}=U_{R}-U_{L}-A_{k}^{-1}S_{k}(H_{R}-H_{L}),$ (63)
have to be solved to compute $\alpha_{i}$, $i=1,\dots,6$ such that (28) is satisfied. Here, $A_{k}=A_{k}(U_{L},U_{R})$, $S_{k}=S_{k}(U_{L},U_{R})$, $k=1,2$ represent a Roe linearization in the $x$ or $y$ direction, and
$R=\left[\begin{array}[]{c|c|c}\vec{R}_{1}&\dots&\vec{R}_{6}\end{array}\right],\quad\vec{\alpha}=\left[\begin{array}[]{c}\alpha_{1}\\ \dots\\ \alpha_{6}\end{array}\right].$
Let us suppose that $k=1$. In this case, $\vec{R}_{i}$, $i=1,\dots,6$ are given by (60). It can be checked that the system can be solved as follows:
@@
    $R=\left[\begin{array}[]{ccc}1&\dots&1\\ \lambda_{1}&\dots&\lambda_{4}\\ \mu_{1}&\dots&\mu_{4}\\ \mu_{1}\lambda_{1}&\dots&\mu_{4}\lambda_{4}\end{array}\right],\quad A=\left[\begin{array}[]{cccccc}0&1&0&0\\ \bar{c}_{1}^{2}-\bar{u}_{1,1}^{2}&2\bar{u}_{1,1}&\bar{c}_{1}^{2}&0\\ 0&0&0&1\\ r\bar{c}_{2}^{2}&0&\bar{c}_{2}^{2}-\bar{u}_{2,1}^{2}&2\bar{u}_{2,1}\end{array}\right],\quad S=\left[\begin{array}[]{c}0\\ g\bar{h}_{1}\\ 0\\ g\bar{h}_{2}\end{array}\right].$
@@
In other words, the system to be solved in the 2D case reduce to those arising in 1D problems. Observe that, following this algorithm, the cases in which eigenvalues $\lambda_{i}$, $i=5,6$, vanish and thus $A_{k}$ cannot be inverted, the coefficients $\alpha_{j}$ can be still computed.
@@ 6 Numerical solutions
In this section, we present some numerical results for the 1D and 2D two-layer shallow water equations.
Two different WENO methods have been implemented and tested:
  • WENO methods with upwind splitting and Strategy 1 in Section 3.1 for the treatment of the source terms: these family of schemes will be called Methods 1.
  • WENO methods with LF splitting and Strategy 2 in Section 3.2 for the treatment of the source terms : these family of schemes will be called Methods 2.
@@
  • the approach in [30] is followed to deal with the episodes of loss of hyperbolicity;
@@
  • the free boundary conditions are imposed excepted in the steady-state solutions and accuracy test.
@@
In many numerical tests the results are comparable and, when this is the case, we only show the results obtained with Methods 2: we will only show results obtained with schemes of both families if there are significant differences or to compare them.

### v2

@@ 4.1 Homogeneous problems
Let us assume again that Roe linearizations $A_{i,\Psi}(U,V)$ of $A_{i}(U)$, $i=1,2$ are available for the selected family of paths $\Psi$. We consider uniform Cartesian meshes with points $(x_{i},y_{j})$ with steps $\Delta x$ and $\Delta y$ in the $x$ and $y$ direction. WENO methods can be extended dimension by dimension:
@@
$D^{\pm}_{i,l;j}$ $=$ $A^{\pm}_{1;i,l;j}(U_{l,j}-U_{i,j}),\quad l=i-k,\dots,i+k,$ (68)
$D^{\pm}_{i;j,l}$ $=$ $A^{\pm}_{2;i;j,l}(U_{i,l}-U_{i,j}),\quad l=j-k,\dots,j+k,$ (69)
where
$A_{1;i,l;j}=A_{1,\Psi}(U_{i,j},U_{l,j}),\quad A_{2;i;j,l}=A_{2,\Psi}(U_{i,j},U_{i,l})$
and the super-indices $\pm$ represent their splitting matrices: both the Upwind and the LF splittings can be readily extended to 2D.
Proposition 1 can be easily extended to prove that (63) is a numerical method of order $p=2k+1$ provided that Property (38) is satisfied for $A_{1}$ and $A_{2}$.
@@ 4.2 Problems with source terms
where $H(x,y)$ is again a known function. Strategy 1 described in Section 3 can be readily extended to problem (70): it is enough to redefine the fluctuations as follows:
$D^{\pm}_{i,l;j}$ $=$ $A^{\pm}_{1;i,l;j}(U_{l,j}-U_{i,j})-S^{\pm}_{1;i,l;j}(H(x_{l},y_{j})-H(x_{i},y_{j})),$ (71)
$D^{\pm}_{i;j,l}$ $=$ $A^{\pm}_{2;i;j,l}(U_{i,l}-U_{i,j})-S^{\pm}_{2;i;j,l}(H(x_{i},y_{l})-H(x_{i},y_{j})),$ (72)
where $A_{1;i,k;j}$, $S_{1;i,k;j}$, represent respectively the intermediate matrix and source term given by the Roe linearization for the states $U_{i,j}$ and $U_{l,j}$ and $A_{2;i;j,l}$, $S_{2;i;j,l}$ the corresponding to the states $U_{i,j}$, $U_{i,l}$.
It can be shown as in Proposition 2 that the numerical method is well-balanced for stationary solutions that satisfy that given $(x_{L},y),(x_{R},y),(x,y_{D}),(x,y_{U})\in\mathbb{R}^{2}$ one has
$A_{1,\widetilde{\Psi}}(U^{*}(x_{L},y),U^{*}(x_{R},y))(U^{*}(x_{R},y)-U^{*}(x_{L},y))$ (73)
$\qquad\qquad=S_{1,\widetilde{\Psi}}(U^{*}(x_{L},y),U^{*}(x_{R},y))(H(x_{R},y)-H(x_{L},y)),$
$A_{2,\widetilde{\Psi}}(U^{*}(x,y_{D}),U^{*}(x,y_{U}))(U^{*}(x,y_{U})-U^{*}(x,y_{D}))$ (74)
$\qquad\qquad=S_{2,\widetilde{\Psi}}(U^{*}(x,y_{D}),U^{*}(x,y_{U}))(H(x,y_{U})-H(x,y_{D}))$
These equalities are satisfied if the functions
@@
$s\in[0,1]\to\Psi_{U}\left(s;\left[\begin{array}[]{c}U^{*}(x,y_{D})\\ H(x,y_{D})\end{array}\right];\left[\begin{array}[]{c}U^{*}(x,y_{U})\\ H(x,y_{U})\end{array}\right]\right),$ (76)
@@
Strategy 2 can be extended easily to 2D problems to design numerical methods that preserve a given family of $m$-parameter stationary solutions
@@
with $m<N$, assuming again that, given $\bar{x},\bar{y},\bar{U}$, there exists a unique choice of the parameters such that
$CU^{*}_{i}(\bar{x},\bar{y})=C\bar{U},$
Once an element of the family $U^{*}_{i,j}$ has been determined, the numerical solution is written as follows:
@@ 5 Application to the two-layer shallow water model
In this Section, we apply the proposed well-balanced schemes in Section 4 to 2D two-layer shallow water system that governs the flow of two superposed layers of immiscible fluids with different constant densities:
@@
$U=\left[\begin{array}[]{c}h_{1},h_{1}u_{1,1},h_{1}u_{1,2},h_{2},h_{2}u_{2,1},h_{2}u_{2,2}\end{array}\right]^{T},$
$A_{1}(U)$ $=\left[\begin{array}[]{cccccc}0&1&0&0&0&0\\ c_{1}^{2}-u_{1,1}^{2}&2u_{1,1}&0&c_{1}^{2}&0&0\\ -u_{1,1}u_{1,2}&u_{1,2}&u_{1,1}&0&0&0\\ 0&0&0&0&1&0\\ rc_{2}^{2}&0&0&c_{2}^{2}-u_{2,1}^{2}&2u_{2,1}&0\\ 0&0&0&-u_{2,1}u_{2,2}&u_{2,2}&u_{2,1}\end{array}\right],\quad S_{1}(U)=\left[\begin{array}[]{c}0\\ gh_{1}\\ gh_{1}\\ 0\\ 0\\ 0\end{array}\right],$ (85)
$A_{2}(U)$ $=\left[\begin{array}[]{cccccc}0&0&1&0&0&0\\ -u_{1,1}u_{1,2}&u_{1,2}&u_{1,1}&0&0&0\\ c_{1}^{2}-u_{1,2}^{2}&0&2u_{1,2}&c_{1}^{2}&0&0\\ 0&0&0&0&0&1\\ 0&0&0&-u_{2,1}u_{2,2}&u_{2,2}&u_{2,1}\\ rc_{2}^{2}&0&0&c_{2}^{2}-u_{2,2}^{2}&0&2u_{2,2}\end{array}\right],\quad S_{2}(U)=\left[\begin{array}[]{c}0\\ 0\\ 0\\ 0\\ gh_{2}\\ gh_{2}\end{array}\right]$
@@
where now $u_{k}$, $h_{k}$, $k=1,2$ are respectively the velocity and thickness of the layers. The family of straight segments will be considered here to compute nonconservative products. The following Roe linearization corresponding to the family of straight segments is available (see [28]):
$A(U_{L},U_{R})=\left[\begin{array}[]{ccccc}0&1&0&0\\ -\bar{u}_{1}^{2}+\bar{c}_{1}^{2}&2\bar{u}_{1}&\bar{c}_{1}^{2}&0\\ 0&0&0&1\\ r\bar{c}_{2}^{2}&0&-\bar{u}_{2}^{2}+\bar{c}_{2}^{2}&2\bar{u}_{2}\end{array}\right],\quad S(U_{L},U_{R})=\left[\begin{array}[]{c}0\\ g\bar{h}_{1}\\ 0\\ g\bar{h}_{2}\end{array}\right],$
@@
This Roe matrix and its natural extension to 2D will be used to implement WENO methods. Steady-states solutions that correspond to water-at-rest equilibria constitute a 2-parameter family:
$u_{1,1}=u_{1,2}=u_{2,1}=u_{2,2}\equiv 0,\;h_{1}=c_{1},\;h_{2}=-Z+c_{2},$
where $c_{1}$, $c_{2}$ are arbitrary parameters ($c_{1}>0$, $c_{2}>\max(Z)$). We will focus here on methods that preserve this family. Both Strategies 1 and 2 described in Section 3 for 1D problems and 4.2 for 2D problems can be followed to design numerical methods that preserve water-at-rest solutions. In effect, for Strategy 1, the equalities (44), (73), (74) can be easily checked for these stationary solutions: the equality of the curves given by (75) and (77) or those given by (76) and (78) can be easily checked if the family of straight segments is chosen: it derives from the linear nature of the relationships between variables that characterize water-at-rest solutions.
For Strategy 2, observe that, given $\bar{x},\bar{y}$ and a state $\bar{U}=[\bar{h}_{1},\bar{h}_{1}\bar{u}_{1,1},\bar{h}_{1}\bar{u}_{1,2},\bar{h}_{2},\bar{h}_{2}\bar{u}_{2,1},\bar{h}_{2}\bar{u}_{2,2}]^{T}$, there exists a unique water-at-rest stationary solution such as
$h_{1}^{*}(\bar{x},\bar{y})=\bar{h}_{1},\quad h_{2}^{*}(\bar{x},\bar{y})=\bar{h}_{2},$
which is the one given by:
$h_{1}^{*}(x,y)=\bar{h}_{1},\quad\bar{h}^{*}_{2}(x,y)=-Z(x,y)+Z(\bar{x},\bar{y})+\bar{h}_{2},\quad u^{*}_{1,1}=u^{*}_{1,2}=u^{*}_{2,1}=u^{*}_{2,2}=0,$
i.e. in this case
$C=\left[\begin{array}[]{cccccc}1&0&0&0&0&0\\ 0&0&0&1&0&0\end{array}\right],$
so that the stationary solution $U^{*}_{i,j}$ used to implement the method is given by
$U^{*}_{i,j}(x,y)=\left[h_{1;i,j},0,0,-Z(x,y)+Z(x_{i},y_{j})+h_{2;i,j},0,0\right]^{T}.$ (89)
When the Upwind splitting scheme is used, systems of the form
$R\cdot\vec{\alpha}=U_{R}-U_{L}-A_{l}^{-1}S_{k}(H_{R}-H_{L}),$
have to be solved to compute $\alpha_{i}$, $i=1,\dots,6$ such that (49) is satisfied. Here, $A_{l}=A_{l}(U_{L},U_{R})$, $S_{l}=S_{l}(U_{L},U_{R})$, $l=1,2$ represent the Roe linearization in the $x$ or $y$ direction, and
$R=\left[\begin{array}[]{c|c|c}\vec{R}_{1}&\dots&\vec{R}_{6}\end{array}\right],\quad\vec{\alpha}=\left[\begin{array}[]{c}\alpha_{1}\\ \vdots\\ \alpha_{6}\end{array}\right].$
Let us suppose that $l=1$. In this case, $\vec{R}_{i}$, $i=1,\dots,6$ are given by (86). It can be checked that the system can be solved as follows:
@@
    $R=\left[\begin{array}[]{ccc}1,&\dots,&1\\ \lambda_{1},&\dots,&\lambda_{4}\\ \mu_{1},&\dots,&\mu_{4}\\ \mu_{1}\lambda_{1},&\dots,&\mu_{4}\lambda_{4}\end{array}\right],\quad A=\left[\begin{array}[]{cccccc}0&1&0&0\\ \bar{c}_{1}^{2}-\bar{u}_{1,1}^{2}&2\bar{u}_{1,1}&\bar{c}_{1}^{2}&0\\ 0&0&0&1\\ r\bar{c}_{2}^{2}&0&\bar{c}_{2}^{2}-\bar{u}_{2,1}^{2}&2\bar{u}_{2,1}\end{array}\right],\quad S=\left[\begin{array}[]{c}0\\ g\bar{h}_{1}\\ 0\\ g\bar{h}_{2}\end{array}\right].$
@@
In other words, the system to be solved in the 2D case reduces to those arising in 1D problems. Observe that, following this algorithm, the cases in which eigenvalues $\lambda_{i}$, $i=5,6$, vanish, and thus $A_{k}$ cannot be inverted, the coefficients $\alpha_{j}$ can be still computed.
@@ 6 Numerical solutions
In this section, we present some numerical results for 1D coupled Burgers, the 1D and 2D two-layer shallow water equations.
Inheriting the notation from Section 3.3, two different WENO methods have been implemented and tested:
  • WENO methods with Upwind splitting and Strategy 1 in Section 3.2.1 for the treatment of the source terms: these family of schemes will be called Method 1.
  • WENO methods with LF splitting and Strategy 2 in Section 3.2.2 for the treatment of the source terms: these family of schemes will be called Method 2.
@@
  • the free boundary conditions are imposed except in the steady-state solutions and accuracy test.
@@
In many numerical tests the results are comparable and, when this is the case, we only show the results obtained with Method 2.
6.1 1D coupled Burgers Equation
In this section, we consider the coupled Burgers’ system (4) and compare the numerical methods obtained with the methods based on the fluctuations (24) and (25) based respectively on the choice of the family of paths $\Psi_{1}$ given by (5) and $\Psi_{2}$ given by (7).
6.1.1 Smooth solution
The initial condition
$\left[u_{0}(x),v_{0}(x)\right]^{T}=\left[u_{0}(x),v_{0}(x)\right]^{T}=\left[0.25\sin(\pi x),0.25\sin(\pi x)\right]^{T}$ (90)
is first consider in the interval $[0,1]$ and periodic boundary conditions are imposed. Figure 1 shows the numerical solutions obtained at time $t=0.1$ with a mesh of 200 points, when the solution is still smooth. Reference solutions have been computed in a mesh of 6400 points and Table 1 shows the errors and the accuracy of the methods: as it can be seen, both methods seem to converge to the same solution, but only the one consistent with the family (5) achieves the optimal accuracy while the one consistent with the family (7) in only first-order accurate. We note that the high-order accuracy condition (38) for family of paths is then necessary.
Figure 1: Section 6.1.1: smooth solution of the Coupled Burgers’ system. Numerical solutions for $u$ (left) and $v$ (right) at $t=0.1$ obtained with the numerical methods consistent with the families of paths $\Psi_{1}$ and $\Psi_{2}$ using a mesh of 200 points.
Table 1: Accuracy test with different paths, $L^{\infty}$ norm error for the coupled Burgers equations.
N path 1 path 2
$L^{\infty}$ error Order $L^{\infty}$ error Order
25 2.13e-6 5.76e-4
50 5.01e-8 5.41 2.86e-4 1.01
100 1.29e-9 5.28 1.42e-4 1.01
200 3.65e-11 5.14 6.98e-5 1.02
400 9.95e-13 5.20 3.38e-5 1.05
6.1.2 Shock waves
The goal of this test is to show that numerical methods based on different families of paths give different results in presence of discontinuous waves. We consider the Riemann problem given by the initial condition
$\left[u_{0}(x),v_{0}(x)\right]^{T}=\begin{cases}\left[2,2\right]^{T},&\text{if}\ x\leq 0.5,\\ \left[1,1\right]^{T},&\text{}\text{otherwise}.\end{cases}$ (91)
The analytical solution is used as the reference solution. If the family of straight segments $\Psi_{1}$ is chosen, the solution is an isolated shock wave traveling at speed 3. After numerically simulating the solution up to $t=0.1$ using 200 uniform grids, the results are presented in Figure 2. As it can be seen, in this particular case the numerical method consistent with the family of paths (5) captures the correct solution, while the one consistent with the family (7) gives a solution with a stationary contact discontinuity at $x=0.5$ and a different shock wave traveling at the same speed.
Figure 2: Section 6.1.2: Riemann problem for the coupled Burgers’ system. Numerical solutions for $u$ (left) and $v$ (right) at $t=0.1$ are obtained consistent with the families of paths $\Psi_{1}$ and $\Psi_{2}$ using a mesh of 200 points.